A buffer solution is 0.371 M in HF and 0.226 M in KF . If Ka for HF is 7.2×10-4, what is the pH of this buffer solution?
Answer #1

CHEJ = 0.3714 [ki] = 0.2264 ka f HF = 7.24104 pka = -log (7.2010) = 4- log (7.2) = 4 - 0.8573 = 3.1426 Apply kunder sons argu